OpenStudy (anonymous):

$216.67-25%=$216.42 Correct Or Not ?

OpenStudy (anonymous):

216.67-25%=216.42 I Put It In On My Calculator And Got This As My Answer But Hold On

OpenStudy (lgbasallote):

why did you subtract 25%??

OpenStudy (anonymous):

Because My Question Ask If The Original Cost Was $216.67 And You Buy Them For 25% Off How Much Is The Cost Of The Boots ?

OpenStudy (anonymous):

\[*\Gamma \Pi \upsilon \delta * \zeta \gamma \xi \epsilon * 4 * \mathbb{R} \epsilon \alpha \zeta *\]

OpenStudy (anonymous):

You Are Pathetic Please Move

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

That's incorrect

OpenStudy (anonymous):

You would divide $216.67 by four, and take that number and subtract it from the original price then you'll get your answer

OpenStudy (anonymous):

$162.50 is the answer!

OpenStudy (anonymous):

The Original Answer Is 216.67

OpenStudy (anonymous):

Well then what's the question How much you get off?

OpenStudy (anonymous):

OHHH WAIT

OpenStudy (anonymous):

It's 270.84

OpenStudy (anonymous):

No If The Original Price Was 216.67 And You Get 25% Off How Much Are They Now ?

OpenStudy (anonymous):

@hartnn Is She Correct Im Lost

OpenStudy (turingtest):

No, she is not correct. @vivalakoda please walk us through how you got your answer.

OpenStudy (anonymous):

She Is Correct Im Sorry Thanks Alot Viva : )

OpenStudy (anonymous):

Well if the original price is 216.67, you first need to find out how much 25% is So divide the original price by 4, this gives you a quarter of the price (54.17) Then since you're taking 25% away, subtract 54.17 from the original price and thus gives us: $162.50.

OpenStudy (anonymous):

(:

OpenStudy (turingtest):

thanks please answer all questions with at least this level of explanation. Better yet, don't give the final answer at all, make the asker do some work themselves!

OpenStudy (anonymous):

I Did Do Work My Self To Thanks ,

OpenStudy (anonymous):

Okay o: I thought this was explaining : "You would divide $216.67 by four, and take that number and subtract it from the original price then you'll get your answer" but okay

OpenStudy (anonymous):

It Is Thanks Girly ,

OpenStudy (anonymous):

you're welcome! :)